OpenStudy (anonymous):

Can someone pls help I have to redo this problem because I got it wrong. 2 - (3x + 1 ) + 2x = 5

jimthompson5910 (jim_thompson5910):

2 - (3x + 1 ) + 2x = 5 2 - 3x - 1 + 2x = 5 -x + 1 = 5 x = ???

OpenStudy (anonymous):

4?

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

-x = 4 x = ???

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

:)

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome